# Simulating thermodynamic properties of dinuclear metal complexes using Variational Quantum Algorithms

### v1

@@ I Introduction
Recently, advancements in quantum processing units (QPUs) have led to the potential for ”quantum supremacy,” where a QPU performs tasks considered impossible for classical computers King et al. (2024). Part of the efforts to harness the capabilities of current quantum devices have resulted in the development of various Quantum Algorithms with a wide range of applications Montanaro (2015). In this regard, the intersection between classical computing and quantum computing gave rise to a class of algorithms that challenge the conventional limits of computing: Quantum Variational Algorithms (VQAs) Cerezo et al. (2021); McClean et al. (2016). These algorithms have emerged as powerful tools for exploring and understanding complex quantum systems, providing valuable insights into their physical behavior and properties.
Based on the theories embedded in the development of VQAs, some innovative approaches to simulating quantum states have been proposed Verdon et al. (2019); Grimsley et al. (2018); Parrish et al. (2019); Tilly et al. (2021); Cerezo et al. (2020); Wu and Hsieh (2018); Li et al. (2019); Zeng et al. (2020); Medina et al. (2023); Endo et al. (2018); Sun et al. (2021); Francis et al. (2020); Kandala et al. (2017); Luo et al. (2024). These algorithms use a quantum circuit and classical optimization routines to simulate the quantum states of a system, given its Hamiltonian. By leveraging the principles of quantum mechanics, VQAs have the potential to revolutionize various fields, such as chemistry and material science Cerezo et al. (2021); Parrish et al. (2019); Grimsley et al. (2018); Li et al. (2019).
@@ II Variational Quantum Algorithms for Simulating Quantum States at Finite Temperature
The complexity and details of the structure of a Quantum Variational Algorithm (VQA) vary depending on the nature of the problem to be solved Cerezo et al. (2021). However, most of these algorithms share fundamental elements such as a cost function (or loss function), a parameterized ansatz, and an optimization loop that optimizes the ansatz Cerezo et al. (2021); Verdon et al. (2019); Tilly et al. (2021). In this context, when facing a specific problem, it is crucial to establish a cost function $\mathrm{C}$ that, based on some variational principle, encodes the solution Verdon et al. (2019); Cerezo et al. (2021). Thus, one can propose an ansatz structured as a quantum operation that depends on a set of parameters $\theta$. The task, therefore, becomes to optimize the ansatz with the aim of finding the parameters that minimize the associated cost.
The structure of the cost function and the details of the ansatz depend on the specific problem being addressed. However, in many cases, the cost is a function of a set of observables $\{O_{k}\}$, a set of input states $\{\rho_{k}\}$, and a parameterized unitary operator, denoted as $U(\theta)$ Cerezo et al. (2021). Thus, this cost function is usually expressed as
$C(\theta)=\sum_{k}f_{k}\left(tr[O_{k}U(\theta)\rho_{k}U^{\dagger}(\theta)]\right)\leavevmode\nobreak\ .$ (1)
@@
VQT is a quantum algorithm that combines machine learning techniques to reproduce thermal states in quantum systems based on a specific Hamiltonian Verdon et al. (2019). This fact is accomplished by approximating the desired thermal state through a quantum-probabilistic ansatz $\widehat{\rho}_{\theta\phi}$, parameterized by the variables ${\theta,\phi}$ Verdon et al. (2019). Thus, the goal of VQT is to tune these parameters in order to minimize the cost $\mathcal{L}_{\theta,\phi}$ given by the relative free energy of the ansatz, concerning a target Hamiltonian $\widehat{H}$:
@@
where $\beta=\dfrac{1}{T}$ is a target inverse temperature and $S(\widehat{\rho}_{\theta,\phi})$ is the von Neumann entropy of the ansatz. This cost function is based on relative entropy, which is also known as the Kullback-Leibler divergence Murphy (2012), which measures the distance between the ansatz and the desired thermal state. This last is denoted as $\widehat{\sigma}_{\beta}$, which is calculated by taking $\dfrac{1}{Z}_{\beta}e^{-\beta\widehat{H}}$, where ${Z}_{\beta}$ represents the canonical partition function of the system at temperature $T$. Thus, by minimizing this cost function, which is achieved by adjusting the ansatz parameters, we can accurately approximate the thermal state of the system.
In this work, we used a variation of the VQT protocol to generate thermal states for a two-qubit system, defined as a dinuclear metal complex with a $d^{9}$ electronic configuration ($s=1/2$) Cruz et al. (2016); Cruz and Anka (2020); Cruz et al. (2023, 2022). Figure 1 depicts a schematic representation of the used protocol. The algorithm begins with an input that includes the Hamiltonian model ($\hat{H}$), the desired reciprocal temperature ($1/T$), and the parameterized probability distribution ($p_{\theta}$). The first step is to define a mixed state $\rho_{\theta}$ using the probability distribution, generating individual pure states. These pure states are then passed through a parameterized quantum circuit $U(\phi)$, which is made up of rotational $RX,\leavevmode\nobreak\ RY,\leavevmode\nobreak\ RZ$ gates and coupling $X$ gates between the two qubits.From the results of this circuit, the expectation value of the Hamiltonian about pure states is calculated. Repeating this process several times and averaging these values gives us the expected value of the Hamiltonian concerning the operator $U(\phi)\rho_{\theta}U^{\dagger}(\phi)$. Due to the invariance of the von Neumann entropy $S_{\theta}$ under unitary transformations, it can be computed in a classical manner using the distribution parameters $\theta_{i}$. After obtaining $\langle\hat{H}\rangle_{\theta,\phi}$ and $S_{\theta}$, we calculate the cost using Eq. (2) and update the parameters until the relative free energy value is minimized. At the end of this hybrid loop, we can use the optimal parameters $\theta,\phi$ to construct the density matrix of our desired mixed state.
@@
Figure 1: Schematic representation of the Variational Quantum Thermalizer based Algorithm: Given a Hamiltonian $\hat{H}$, a target inverse temperature, and a parameterized probability distribution $p_{\theta}$, the VQT starts with a mixed state $\rho_{\theta}$ described by the probability distribution. We generate instances of this probability distribution to obtain individual pure states, which are then passed through a parameterized quantum circuit $U(\phi)$, composed of rotational $RX,\leavevmode\nobreak\ RY,\leavevmode\nobreak\ RZ$ and coupling $X$ between the qubits (in our case, 2 qubits). From the results of this circuit, the expectation value of the Hamiltonian about pure states is calculated. Repeating this process several times and averaging these values gives us the expected value of the Hamiltonian concerning $U(\phi)\rho_{\theta}U^{\dagger}(\phi)$. Due to the invariance of the von Neumann entropy $S_{\theta}$ under unitary transformations, it is calculated classically using the parameters $\theta_{i}$ of the distribution. Once $\langle\hat{H}\rangle_{\theta,\phi}$ and $S_{\theta}$ were obtained, we calculate the cost given by Eq. (2), and the parameters are updated until the relative free energy value is minimum. At the end of this hybrid loop, using the optimal parameters $\theta,\phi$, we can construct the density matrix of our target mixed state.
@@
To implement the quantum algorithm, we used Pennylane Bergholm et al. (2022), an open-source Python library that offers a flexible platform for simulating materials using quantum circuit simulations, containing advanced methods for VQAs, preparation of thermal states, and evaluation of molecular spectra Bassman et al. (2021). For the classical optimization problem, we utilized the Constrained Optimization BY Linear Approximation (COBYLA) method Powell (1994). COBYLA is a numerical optimization technique designed for derivative-free constrained optimization problems where the derivative of the objective function is not known. It involves an iterative approach for derivative-free constrained optimization problems that have been advanced and refined over the years and is widely used in quantum-classical hybrid algorithms Cheng et al. (2024); Singh et al. (2023); Pellow-Jarman et al. (2021).
In the following, we present the assessed thermodynamic properties of the material under investigation and discuss the implications of using quantum computing for future material engineering. We demonstrate the accuracy of the approach discussed in this section in predicting the thermodynamic properties of the studied material.
@@ III Thermodynamic Properties of a Dinuclear Metal Complex
$=$ $J\left[\begin{matrix}1&0&0&0\\ 0&-1&2&0\\ 0&2&-1&0\\ 0&0&0&1\end{matrix}\right]\leavevmode\nobreak\ ,$
@@
$\rho$ $=\dfrac{1}{2}\left[\begin{matrix}2\varrho_{2}&0&0&\\ 0&{\varrho_{1}+\varrho_{3}}&{\varrho_{3}-\varrho_{1}}&0\\ 0&{\varrho_{3}-\varrho_{1}}&{\varrho_{1}+\varrho_{3}}&0\\ 0&0&0&2\varrho_{4}\end{matrix}\right].$ (13)
@@ IV Simulating Thermodynamic Properties Through Quantum Computing
This section will present the results obtained using the variational algorithm presented in section II for the dinuclear spin $1/2$ metal complex described. Figure 2 depicts the tomography of the system Hamiltonian in the computational basis, as defined by Eq. (7). As can be seen, the matrix entries are parameterized by the system’s magnetic coupling constant $J$.
Figure 2: Simulated Hamiltonian matrix, parametrized by the magnetic coupling constant $J$ in the computational basis.
Figure 3: Thermal states obtained in the simulation for different temperature values: (a) thermal state at $T=5\times 10^{-4}\leavevmode\nobreak\ K$; (b) $T=10\leavevmode\nobreak\ K$; (c) $T=20\leavevmode\nobreak\ K$; (d) $T=350\leavevmode\nobreak\ K$. All these thermal states were reproduced after $1600$ simulation steps, with $J/k_{B}=1$ K, where $J$ and $k_{B}$ are the magnetic coupling and the Boltzmann constant, respectively.
As can be seen, considering $J/k_{B}=1$, Figure 3 (a) shows that at very low temperatures ($T=5\times 10^{-4}\leavevmode\nobreak\ K$), the system exists in the pure singlet ground state $|\varrho_{4}\rangle=\frac{1}{\sqrt{2}}\left(|01\rangle{-}|10\rangle\right)$, which is a maximally entangled mixed state. As we increase the temperature to $10\leavevmode\nobreak\ K$ in Figure 3 (b), the density matrix approaches a two-qubit X-shaped state Sarandy (2009), which is characterized by the terms of the secondary diagonal of the matrix. Upon reaching $20\leavevmode\nobreak\ K$ in Figure 3 (c), the populations increase while the coherence terms decrease. Finally, when the temperature reaches a high enough value of $T=350\leavevmode\nobreak\ K$, the system tends towards a classical probability distribution, which is characterized by the density matrix in its diagonal form, as shown in Figure 3 (d). This evolution from an entangled mixed state to a classical probability distribution highlights the transition from quantum to classical behavior as temperature increases. The changes in the density matrix elements reflect the diminishing quantum correlations and the emergence of classical-like features in the system as thermal fluctuations dominate Schlosshauer (2019).
@@
As presented in section III, the simulated thermal state provides valuable insights into the thermodynamic properties of the system, offering a deeper understanding of the influence of the macroscopic properties of the system in its behavior at the quantum level Cruz et al. (2016). We evaluate the thermal states reproduced by the variational algorithm for a set of temperature values, from $T=5\times 10^{-4}\leavevmode\nobreak\ K$ to $350\leavevmode\nobreak\ K$.
@@
Analyzing the magnetic susceptibility of magnetic systems is a crucial step in comprehending how magnetic properties react to changes in temperature and magnetic coupling ($J$) Reis (2013). From the simulated thermal state, we can obtain the molar magnetic susceptibility, Eq. (16), as a function of the temperature for different magnetic couplings. Figure 4 shows the product $T\chi(T)$ as a function of the temperature. The behavior of this function is consistent with Curie’s Law in the paramagnetic region, where the magnetic orientation of the spins becomes less ordered due to thermal agitation, resulting in the loss of the intrinsic magnetic properties of the system in the paramagnetic transition Reis (2013).
@@
Figure 4: Dimentionless molar magnetic susceptibility, Eq. (16), times the temperature for different coupling values $J$. The data were obtained through thermal states reproduced by the variational quantum algorithm presented in section II considering temperature values in the range of $5\times 10^{-5}\leavevmode\nobreak\ K$ to $350\leavevmode\nobreak\ K$. Dashed line curves represent an analytical simulation of the Bleaney-Bowers equation presented in Eq. (16).
@@
Figure 5: (a) Entropy of the Heisenberg dimer as a temperature function for different coupling values $J$. The data were obtained through thermal states reproduced by VQT (Variational Quantum Thermalizer Algorithm) considering temperature values in the range of $5\times 10^{-5}\leavevmode\nobreak\ K$ to $350\leavevmode\nobreak\ K$. (b) Magnetic specific heat as a function of temperature for different values of magnetic coupling $J$. The data points were derived from entropy data. The curves with dashed lines refer to the theoretical model for the simulated Hamiltonian, Eq. (7).

### v2

@@ I Introduction
Recently, advancements in quantum processing units (QPUs) have led to the potential for ”quantum supremacy,” where a QPU performs tasks considered impossible for classical computers King et al. (2024). Part of the efforts to harness the capabilities of current quantum devices have resulted in the development of various Quantum Algorithms with a wide range of applications Montanaro (2015). In this regard, the intersection between classical computing and quantum computing gave rise to a class of algorithms that challenge the conventional limits of computing: Variational Quantum Algorithms (VQAs) Cerezo et al. (2021); McClean et al. (2016). These algorithms have emerged as powerful tools for exploring and understanding complex quantum systems, providing valuable insights into their physical behavior and properties.
Based on the theories embedded in the development of VQAs, some innovative approaches to simulating quantum states have been proposed Verdon et al. (2019); Grimsley et al. (2018); Parrish et al. (2019); Tilly et al. (2021); Cerezo et al. (2020); Wu and Hsieh (2018); Li et al. (2019); Zeng et al. (2020); Medina et al. (2023); Endo et al. (2018); Sun et al. (2021); Francis et al. (2020); Kandala et al. (2017); Luo et al. (2024); Consiglio et al. (2023); Holmes et al. (2022). These algorithms use a quantum circuit and classical optimization routines to simulate the quantum states of a system, given its Hamiltonian. By leveraging the principles of quantum mechanics, VQAs have the potential to revolutionize various fields, such as chemistry and material science Cerezo et al. (2021); Parrish et al. (2019); Grimsley et al. (2018); Li et al. (2019).
@@ II Variational Quantum Algorithms for Simulating Quantum States at Finite Temperature
The complexity and details of the structure of a Quantum Variational Algorithm (VQA) vary depending on the nature of the problem to be solved Cerezo et al. (2021). However, most of these algorithms share fundamental elements such as a cost function (or loss function), a parameterized ansatz, and an optimization loop that optimizes the ansatz Cerezo et al. (2021); Verdon et al. (2019); Tilly et al. (2021). In this context, when facing a specific problem, it is crucial to establish a cost function $\mathrm{C}$ that, based on some variational principle, encodes the solution Verdon et al. (2019); Cerezo et al. (2021). Thus, one can propose an ansatz structured as a quantum operation that depends on a set of parameters $\phi$. The ansatz is an important part of variational algorithms because it sets up the initial quantum state so that it matches the wave function of the circuit: $U(\phi)|{\Psi_{0}}\rangle=|{\Psi(\phi)}\rangle$. Therefore, the task is to optimize the ansatz to find the parameters that minimize the associated cost, following the variational principle Cerezo et al. (2020):
$E_{min}\leq E(\phi)=\langle{\Psi(\phi)}|\mathcal{H}|{\Psi(\phi)}\rangle\leavevmode\nobreak\ ,$ (1)
which guarantees a good approximation of the ground state of the system’s Hamiltonian $\mathcal{H}$. The structure of the cost function and the details of the ansatz depend on the specific problem being addressed. However, in many cases, the cost is a function of a set of observables $\{O_{k}\}$, a set of input states $\{\rho_{k}\}$, and the parameterized unitary operator $U(\phi)$ Cerezo et al. (2021). Thus, this cost function is usually expressed as
$C(\phi)=\sum_{k}f_{k}\left(tr[O_{k}U(\phi)\rho_{k}U^{\dagger}(\phi)]\right)\leavevmode\nobreak\ .$ (2)
@@
VQT is a quantum algorithm that combines machine learning techniques to reproduce thermal states in quantum systems based on a specific Hamiltonian Verdon et al. (2019). This fact is accomplished by approximating the desired thermal state $\widehat{\sigma}_{\beta}$, given by $e^{-\beta\widehat{H}}/{Z}_{\beta}$, where ${Z}_{\beta}$ is the canonical partition function of the system at temperature $T$, through a quantum-probabilistic ansatz $\widehat{\rho}_{\theta\phi}$, parameterized by the variables ${\theta,\phi}$ Verdon et al. (2019). Thus, the goal of VQT is to tune these parameters to minimize the cost $\mathcal{L}_{\theta,\phi}$ given by the relative free energy of the ansatz concerning a target Hamiltonian $\widehat{H}$:
@@
where $\beta={1}/{T}$ is a target inverse temperature and $S(\widehat{\rho}_{\theta,\phi})$ is the von Neumann entropy of the ansatz. This cost function, Eq. (3), can be understood in terms of the quantum relative entropy $D(\widehat{\rho}_{\theta,\phi}||\widehat{\sigma}_{\beta})$ Verdon et al. (2019), which is defined as:
$D(\widehat{\rho}_{\theta,\phi}||\widehat{\sigma}_{\beta})$ $=$ $-S(\widehat{\rho}_{\theta,\phi})-\text{tr}(\widehat{\rho}_{\theta,\phi}\log\widehat{\sigma}_{\beta})$ (4)
$=$ $-S(\widehat{\rho}_{\theta,\phi})+\beta\text{tr}(\widehat{\rho}_{\theta\phi}\widehat{H})+\log{{Z}_{\beta}}$
$=$ $\mathcal{L}_{\theta,\phi}+\log{{Z}_{\beta}}\leavevmode\nobreak\ .$
Except for the partition function $Z_{\beta}$, which does not depend on the variational parameters $\theta,\phi$ and therefore does not influence the optimization, our cost function, shown in Eq. (3), is directly defined as the quantum relative entropy, as given by Eq. (4) Verdon et al. (2019). It quantifies the loss of information that results from incorrectly describing a system that is actually in state $\hat{\sigma}_{\beta}$ as being in state $\hat{\rho}_{\theta\phi}$ Murphy (2012); Potts (2019). Thus, by minimizing this cost function, which is achieved by adjusting the ansatz parameters, we can accurately approximate the thermal state of the system.
In this work, we used the VQT protocol Verdon et al. (2019) to generate thermal states for a two-qubit system, defined as a dinuclear metal complex with a $d^{9}$ electronic configuration ($s=1/2$) Cruz et al. (2016); Cruz and Anka (2020); Cruz et al. (2023, 2022). Figure 1 depicts a schematic representation of the used protocol. The algorithm begins with an input that includes the Hamiltonian model ($\hat{H}$), the desired reciprocal temperature ($1/T$), and the parameterized probability distribution ($p_{\theta}$).
@@
Figure 1: Schematic representation of the Variational Quantum Thermalizer based Algorithm: Given a Hamiltonian $\hat{H}$, a target inverse temperature, and a parameterized probability distribution $p_{\theta}$, the VQT starts with a mixed state $\rho_{\theta}$ described by the probability distribution. We generate instances of this probability distribution to obtain individual pure states, which are then passed through a parameterized quantum circuit $U(\phi)$, composed of rotational $RX,\leavevmode\nobreak\ RY,\leavevmode\nobreak\ RZ$ and controlled-X rotation gate between the qubits (in our case, 2 qubits). From the results of this circuit, the expectation value of the Hamiltonian about pure states is calculated. Repeating this process several times and averaging these values gives us the expected value of the Hamiltonian concerning $U(\phi)\rho_{\theta}U^{\dagger}(\phi)$. Due to the invariance of the von Neumann entropy $S_{\theta}$ under unitary transformations, it is calculated classically using the parameters $\theta_{i}$ of the distribution. Once $\langle\hat{H}\rangle_{\theta,\phi}$ and $S_{\theta}$ were obtained, we calculate the cost given by Eq. (3), and the parameters are updated until the relative free energy value is minimum. At the end of this hybrid loop, using the optimal parameters $\theta,\phi$, we can construct the density matrix of our target mixed state.
In this regard, the algorithm’s first step is to define a mixed state $\rho_{\theta_{i}}$, for the two-qubit system in its factored form:
$\rho_{\theta_{i}}=\rho_{\theta_{i}}^{(1)}\otimes\rho_{\theta_{i}}^{(2)}\leavevmode\nobreak\ .$ (5)
For each single-qubit system, we have
${\rho_{\theta_{i}}^{(\alpha)}=p_{i}(\theta_{i})|{0}\rangle\langle{0}|+(1-p_{i}(\theta_{i}))|{1}\rangle\langle{1}|}\leavevmode\nobreak\ ,$ (6)
where we use the probability distribution $p_{i}(\theta_{i})$ in order to generate pure states, with $\alpha=\{1,2\}$. It is worth highlighting that the advantage of using this factored model is that the number of parameters $\theta_{i}$ needed to prepare the state scales linearly rather than exponentially with the number of subsystems Verdon et al. (2019). Thus, the exact number of parameters depends on the configuration of states within each subsystem Verdon et al. (2019). The pure states are then passed through a parameterized quantum circuit $U(\phi)$, which is made up of rotational $RX,\leavevmode\nobreak\ RY,\leavevmode\nobreak\ RZ$ gates and controlled-X rotation gates between the two qubits, as can be seen in Figure 1.
From the results of this circuit, the expectation value of the Hamiltonian about pure states is calculated. After repeating this process several times and averaging these values, the algorithm gives us the expected value of the Hamiltonian concerning the operator $U(\phi)\rho_{\theta}U^{\dagger}(\phi)$. Due to the invariance of the von Neumann entropy $S_{\theta}$ under unitary transformations, it can be computed in a classical manner using the distribution parameters $\theta_{i}$. After obtaining $\langle\hat{H}\rangle_{\theta,\phi}$ and $S_{\theta}$, we calculate the cost using Eq. (3) and update the parameters until the relative free energy value is minimized. At the end of this hybrid loop, we can use the optimal parameters $\theta,\phi$ to construct the density matrix of our desired mixed state.
@@
For the classical optimization problem, we utilized the Constrained Optimization BY Linear Approximation (COBYLA) method Powell (1994). COBYLA is a numerical optimization technique designed for derivative-free constrained optimization problems where the objective function’s derivative is unknown. It involves an iterative approach for derivative-free constrained optimization problems that have been advanced and refined over the years and is widely used in quantum-classical hybrid algorithms Cheng et al. (2024); Singh et al. (2023); Pellow-Jarman et al. (2021).
In this regard, to implement the quantum algorithm in an ideal quantum processor we used Pennylane’s Statevector Simulator Bergholm et al. (2022). Pennylane is an open-source Python library that offers a flexible platform for simulating materials using quantum circuit simulations, containing advanced methods for VQAs, preparation of thermal states, and evaluation of molecular spectra Bassman et al. (2021). Furthermore, to analyze the algorithm’s performance in a more realistic approach, we used the IBM Quantum Experience simulators, qasm_simulator and FakeSheerbrooke in the IBM Qiskit library Javadi-Abhari et al. (2024). The qasm_simulator in Qiskit is a classical backend for simulating quantum circuits described in a shot-based environment. FakeSheerbrooke is a noisy simulator in Qiskit that mimics the noise characteristics of IBM’s Sherbrooke real quantum computer. It allows users to test quantum algorithms in an environment that closely resembles the real Sherbrooke quantum device, including its realistic noise model.
It is important to emphasize that the results presented in this paper were obtained by rebuilding Pennylane’s built-in routine Ceroni (2020) to meet the specifics of the simulation of metal complex systems. In this regard, we implement a routine to handle the variation of the system’s magnetic coupling between the qubits, which was not possible in the original algorithm. This adjustment is fundamental for control and experimentation in our context, as we were able to analyze the thermodynamic properties for a wide range of temperatures and different coupling values ($J$) between the spins. In the following, we present the assessed thermodynamic properties of the material under investigation and discuss the implications of using quantum computing for future material engineering. We demonstrate the accuracy of the approach discussed in this section in predicting the thermodynamic properties of metal complexes.
@@ III Thermodynamic Properties of a Dinuclear Metal Complex
$=$ ${\frac{J}{4}}\left[\begin{matrix}1&0&0&0\\ 0&-1&2&0\\ 0&2&-1&0\\ 0&0&0&1\end{matrix}\right]\leavevmode\nobreak\ ,$
@@
$\rho$ $=\dfrac{1}{2}\left[\begin{matrix}2\varrho_{2}&0&0&{0}\\ 0&{\varrho_{1}+\varrho_{3}}&{\varrho_{3}-\varrho_{1}}&0\\ 0&{\varrho_{3}-\varrho_{1}}&{\varrho_{1}+\varrho_{3}}&0\\ 0&0&0&2\varrho_{4}\end{matrix}\right].$ (9)
@@ IV Simulating Thermodynamic Properties Through Quantum Computing
This section will present the results obtained using the variational algorithm presented in section II for the described dinuclear spin $1/2$ metal complex. Equation (7) defines the form of the system’s Hamiltonian. As can be seen, the matrix entries are parameterized by the system’s magnetic coupling constant $J$. In appendix A we show the Hamiltonian and density matrix tomography obtained from the statevector simulation.
@@
In the following, the Hamiltonian can be used to obtain the thermal state, Eq. (9). The results reveal the probability distributions $\varrho_{n}$, Eqs. (10) and (11), for each temperature value (see the Appendix A for more details). As presented in section III, the simulated thermal state provides valuable insights into the thermodynamic properties of the system, offering a deeper understanding of the influence of the macroscopic properties of the system in its behavior at the quantum level Cruz et al. (2016).
We evaluate the thermal states reproduced by the variational algorithm (see section II) for a set of temperature values, from $T=5\times 10^{-4}\leavevmode\nobreak\ K$ to $350\leavevmode\nobreak\ K$. Figure 2 shows the performance of minimizing the cost function, given by Eq. (3), during the simulation for high and low-temperature values. In all cases analyzed, the model was trained for 400 randomly chosen initial parameter sets with $J/k_{B}=1\leavevmode\nobreak\ K$, and the cost function converges to a minimum value after approximately 200 iterations.
Figure 2: Performance of the optimization of the cost function given by the Eq. 3 for different temperature values: (a) $T=5\times 10^{-4}\leavevmode\nobreak\ K$, (b) $10\leavevmode\nobreak\ K$, (c) $20\leavevmode\nobreak\ K$ and (d) $300\leavevmode\nobreak\ K$. For optimization, we used the Constrained Optimization BY Linear Approximation (COBYLA) method. The model was trained for 400 randomly chosen sets of initial parameters with $J/k_{B}=1\leavevmode\nobreak\ K$.
@@
Analyzing the magnetic susceptibility of magnetic systems is a crucial step in comprehending how magnetic properties react to changes in temperature and magnetic coupling ($J$) Reis (2013). From the simulated thermal state, we can obtain the molar magnetic susceptibility, Eq. (12), as a function of the temperature for different magnetic couplings. Figure 3 shows the product $T\chi(T)$ as a function of the temperature. The behavior of this function is consistent with Curie’s Law in the paramagnetic region, where the magnetic orientation of the spins becomes less ordered due to thermal agitation, resulting in the loss of the intrinsic magnetic properties of the system in the paramagnetic transition Reis (2013). The magnetic coupling values were selected from 1 K to 50 K to present a wide range of magnetic susceptibility, spanning from low temperatures to near room temperature, approaching the paramagnetic region.
@@
Figure 3: Dimentionless molar magnetic susceptibility, Eq. (12), times the temperature for different coupling values $J$. The data were obtained through thermal states reproduced by the variational quantum algorithm presented in section II considering temperature values in the range of $5\times 10^{-5}\leavevmode\nobreak\ K$ to $350\leavevmode\nobreak\ K$. Dashed line curves represent an analytical simulation of the Bleaney-Bowers equation presented in Eq. (12). The magnetic coupling values were selected to ease the visualization of Schottky’s anomaly of the specific heat.
@@
Figure 4: (a) Entropy of the Heisenberg dimer as a temperature function for different coupling values $J$. The data were obtained through thermal states reproduced by VQT (Variational Quantum Thermalizer Algorithm) considering temperature values in the range of $5\times 10^{-5}\leavevmode\nobreak\ K$ to $200\leavevmode\nobreak\ K$. (b) Magnetic specific heat as a function of temperature for different values of magnetic coupling $J$. The data points were derived from entropy data. The curves with dashed lines refer to the theoretical model for the simulated Hamiltonian, Eq. (7).